OpenStudy (bubba):

You have a 3-card deck containing a king, a queen, and a jack. You draw a random card, then put it back and draw a second random card. Calculate the probability that you draw exactly 1 queen.

OpenStudy (anonymous):

1/3 +1/3 - 1/9

OpenStudy (anonymous):

1/3 * 2/3 + 2/3 * 1/3 = 2/6 + 2/6 = 2/3

OpenStudy (anonymous):

Use a ttre diagram to see how it works

OpenStudy (anonymous):

The probability is based on 2 draws one of which can be a queen for a successful outcome. There are two ways for this to happen. Q then not Q or Not Q the Q P(Q) = 1/3 therefore P(not Q) = 2/3. Multiply P(Q) by P(not Q) twice then add the result.

OpenStudy (anonymous):

Ok so now I will ammend my initial response based on my last post 2/9 + 2/9 =4/9 We got there in the end!

OpenStudy (anonymous):

i am afraid gianfranco is right and i am wrong. very sorry. it is queen on first draw not queen on second draw = 1/3*2/3 or not queen on first draw, queen second draw = 2/3*1/3 since they are mutually exclusive "or " means add and you get 4/9